Reduce the following fractions to the simplest form: 
 7 / 28

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

7 / 28 = 7 / (7 × 4)
= 1/4

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Simplest form:
7/28
Divide by 7 in both numerator and denominator,
= 7÷7/28÷7
= 1/4
 
Hope it helps!
Thank you